# MCGAN: Enhancing GAN Training with Regression-Based Generator Loss

### v1

@@ A.4 Discriminability
    where $f^{\prime}$ is the first order derivative of $f$ and increasing due to the convexity of $f$. Consequently, we can choose $c=f^{\prime}(1)$ such that $(D^{\phi^{*}_{\mu,\nu_{\theta}}}(x)-c)(p_{\mu}(x)-p_{\nu_{\theta}}(x))>0$ when $p_{\mu}(x)\neq p_{\nu_{\theta}}(x).$ A more detailed list of $f$-divergence can be found in [44].
Appendix B Architectures, hyperparameters, and training techniques
B.1 BigGAN architecture
Table 5 shows the architectural details.
@@ B.1 BigGAN architecture
B.2 Choice of MC size
In the BigGAN experiment, the MC size we used is 10. While in the StyleGAN experiment, the MC size is only 4. While the MC size we choose is surprisingly small, MCGAN still achieves superior performance on conditional image generation tasks. It might be counter-intuitive that a large MC size will not necessarily result in a better generator. Although a larger MC size will lead to more accurate estimated conditions, it might deteriorate the generalization of the generator network. As explored in [46], large batch size usually leads to a loss in generalization performance. While the small batch method achieves low precision, it tends to have better performance on the test set than the large batch size which is very similar to the bais-variance trade-off. For the same reason, we would not choose a large MC size which will not necessarily lead to better performance, not to mention the enormously increased computational cost it can raise.
B.3 Weight decay
To prevent overfitting of the discriminator, we apply weight decay. It is observed that a high weight decay value can deteriorate the model performance, hence we need to choose the weight decay value carefully. We also found empirically that the value of weight decay is often related to the diversity of the training set. Hence the weight decay we suggest for the discriminator is 0.001 in the CIFAR-10 experiment and 0.005 for the CIFAR-100 experiment, as we have a very limited training sample per class in the CIFAR-100 dataset. When combining weight decay with differentiable augmentation, we chose a low level of weight decay for the discriminator, that is 0.0001 as the differentiable augmentation also improved the diversity of training samples.
As recommended in [11], we also applied weight decay to the generator as it was found that the generator can also have an overfitting problem and that applying weight decay on the generator makes the training more stable. That is probably due to the fact that at the late stage of training, the generator focus on generating images that achieves high logits, which harms the diversity of generated samples. Hence we also set the weight decay of the generator as 0.0001 to avoid the risk of overfitting.
Appendix C Algorithm
The algorithm for the conditional generation task can be found in Algorithm 1.
Input:
     $N$: the number of epochs;
     $N_{D}$: number of discriminator iterations per generator iteration;
     $B\in\mathbb{N}$: batch size;
     $N_{MC}$: the number of Monte Carlo samples;
     $f$: specified function used to compute discriminative loss.
     $C_{(lb,ub)}$: the clamp function with lower bound $lb$ and upper bond $ub$.
Output:
     $(\theta^{*},\psi^{*})$: approximation of the optimal parameters of the generator and discriminator.
1:Initialize the model parameter $(\theta,\psi)$ of the generator $G$ and discriminator $D$. for $n=1:N$ do
       for $n_{d}=1:N_{D}$ do
2:Sample batch $\{(x_{i},y_{i})\}^{B}_{i=1}\sim p_{d}(X,Y)$
3:Generate fake samples given labels $\{(\hat{x}_{i},y_{i})\}^{B}_{i=1}\sim p_{\theta}(X,Y)$
4:Compute discriminative loss:
$\mathcal{L}_{D}(\phi;\mu,\nu_{\theta})\leftarrow\frac{1}{B}\sum_{i=1}^{B}f_{1}(D^{\psi}(x_{i},y_{i}))+\frac{1}{B}\sum_{i=1}^{B}f_{2}(D^{\psi}(\hat{x}_{i},y_{i}));$
            5:Update parameter of discriminator:
$\psi\leftarrow\text{Adam}(\mathcal{L}_{D})$
       end for
6:Sample batch $\{(x_{i},y_{i})\}^{B}_{i=1}\sim p_{d}(X,Y)$
7:For each label $y_{i}$, estimate the conditional expectation
$\hat{\mathbb{E}}_{p_{\theta}}[C_{(lb,ub)}(D^{\psi}(X,y_{i}))|y_{i}]\leftarrow$
${\frac{1}{N_{mc}}\sum_{j=1}^{N_{mc}}C_{(lb,ub)}(D^{\psi}(G^{\theta}(y_{i},z^{(j)}),y_{i}))}$
8:Compute generative loss:
$\mathcal{L}_{G}\leftarrow$
$\frac{1}{B}\sum_{i=1}^{B}\|C_{(lb,ub)}D^{\psi}(x_{i},y_{i}))-\hat{\mathbb{E}}_{p_{\theta}}[C_{(lb,ub)}(D^{\psi}(X,y_{i}))|y_{i}]\|^{2};$
      9:Update parameter of generator:
$\theta\leftarrow\text{Adam}(\mathcal{L}_{G})$
end for
ALGORITHM 1 Algorithm of MCGAN

### v3

@@ Appendix C Proofs
    where $f^{\prime}$ is the first order derivative of $f$ and increasing due to the convexity of $f$. Consequently, we can choose $c=f^{\prime}(1)$ such that $(D^{\phi^{*}_{\mu,\nu_{\theta}}}(x)-c)(p_{\mu}(x)-p_{\nu_{\theta}}(x))>0$ when $p_{\mu}(x)\neq p_{\nu_{\theta}}(x).$ A more detailed list of $f$-divergence can be found in (Nowozin, Cseke, and Tomioka 2016).
Appendix D Evaluation metrics
In this section, we provide detailed introduction to those test metrics used in our numerical experiment.
Evaluation metrics of image generation
Inception Score (Salimans et al. 2016) (IS) is a popular metric to evaluate the variety and distinctness of the generated images. It is given by
$\text{IS}=\exp\{\mathbb{E}_{X\sim\nu_{\theta}}\left[D_{\text{KL}}(P(Y|X)||P(Y))\right]\},$ (25)
where $D_{\text{KL}}$ is the KL-divergence between the conditional class distribution $P(Y|X)$ and marginal class distribution of the $P(Y)=\mathbb{E}[P(Y|\mathbf{X})]$. The conditional class distribution $P(Y|X)$ is computed by InceptionV3 network pre-trained on ImageNet. The higher IS, the better the quality. By the definition, the IS does not consider real images, so cannot measure how well the fake measure induced by the generator is close to the real distribution. Other limitations, as noted in (Barratt and Sharma 2018), are: high sensitivity to small changes in weights of the Inception network, and large variance of scores. To consider both diversity and realism, the following FID and IFID are employed as well.
Fréchet Inception Distance (Heusel et al. 2017) (FID) compares the distributions of Inception embeddings of real and generated images, denoted by $p_{d}$ and $p_{\theta}$ respectively. Under the assumption that the features of images extracted by the function $f$ are of multivariate normal distribution. The FID score of $p_{\theta}$ w.r.t $p_{d}$ is defined as
$\text{FID}(p_{d},p_{\theta})=$ $\|\mathbf{\mu}_{r}-\mathbf{\mu}_{g}\|^{2}$ (26)
$+\text{Tr}(\mathbf{\Sigma}_{r}+\mathbf{\Sigma}_{g}-2(\mathbf{\Sigma}_{r}\mathbf{\Sigma}_{g})^{\frac{1}{2}}),$
$\mathbf{\mu}$ $=\frac{1}{N}\sum^{N}_{i=0}f(x_{i}),$
$\Sigma$ $=\frac{1}{N-1}\sum_{i=0}^{N}(f(x^{(i)})-\mathbb{\mu})(f(x^{(i)})-\mathbb{\mu})^{T}.$
Intra Fréchet Inception Distance (DeVries et al. 2019) (IFID) is used to quantify intra-class diversity. It is defined as the average of conditional FID given every class $y\in\mathcal{Y}$, i.e.,
where
$\text{FID}(p_{d}(y),p_{\theta}(y))=\|\mathbf{\mu}_{r}(y)-\mathbf{\mu}_{g}(y)\|^{2}$
$+\text{Tr}(\mathbf{\Sigma}_{r}(y)+\mathbf{\Sigma}_{g}(y)-2(\mathbf{\Sigma}_{r}(y)\mathbf{\Sigma}_{g}(y))^{\frac{1}{2}}).$
Recognizability is as crucial as realism and diversity in a good image generative model, therefore two classification accuracy are adopted—a weak accuracy (WA) measured by a two-layer convolutional neural network  [^4] https://pytorch.org/tutorials/beginner/blitz/cifar10_tutorial.html and a strong accuracy (SA) by the ResNet-50 (He et al. 2016). Both classifiers are pre-trained on the same training set as for the generative model. The WA discerns subtle differences for better inter-model comparison, while the SA is more accurate for intra-model latent space analysis.
Evaluation metrics of timeseries generation
In the following context, we describe the definition of the test metrics precisely, more detailed discussions can be found in (Liao et al. 2024; Xiao 2023).
  • ABS metric on marginal distribution: The ABS metric is a histogram-based distributional metrics where we compare the empirical density function (epdf) of real data and synthetic data. When talking about epdf, we mean each bin’s raw count divided by the total number of counts and the bin width. For each feature dimension $i\in\{1,\cdots,d\}$, we denote the epdfs of real data and synthetic data as $\hat{df}_{r}^{i}$ and $\hat{df}_{G}^{i}$ respectively. Here the epdfs of synthetic date $\hat{df}_{G}^{i}$ is computed on the bins derived from the histogram of real data. The ABS metric is defined as the absolute difference of those two epdfs averaged over feature dimension, i.e.
    $\frac{1}{d}\sum_{i=1}^{d}|\hat{df}_{r}^{i}-\hat{df}_{G}^{i}|_{1},$
    where $|\hat{df}_{r}^{i}-\hat{df}_{G}^{i}|_{1}$ is computed as the $l_{1}$ distance between the epdfs of real and synthetic data on each bin. Notice that although the ABS metric cannot give a fully point-separating metric on the space of measure, it can still provide a general description of the similarity between two set of data given a reasonable number of bins. Considering the computational cost, we set number of bins to 50 in our implementation.
  • ACF metric on temporal dependency: We use the absolute error of the auto-correlation estimator by real data and synthetic data as the metric to assess the temporal dependency and name it as ACF metric. For each feature dimension $i\in\{1,\ldots,d\}$, we compute the auto-covariance of the $i^{th}$ coordinate of time series data $X$ at lag $\tau$ under real measure and synthetic measure resp, denoted by $\rho_{r}^{i}(\tau)$ and $\rho_{G}^{i}(\tau)$. Then the estimator of the lag-$\tau$ auto-correlation of the real/synthetic data is given by $\frac{\rho_{r}^{i}(\tau)}{\rho_{r}^{i}(0)}$/ $\frac{\rho_{G}^{i}(\tau)}{\rho_{G}^{i}(0)}$. The ACF metric is defined to be the absolute difference of auto-correlation up to lag $\tau$ given as follows:
    $\frac{1}{d\tau}\sum_{k=1}^{\tau}\sum_{i=1}^{d}\left|\frac{\rho_{r}^{i}(k)}{\rho_{r}^{i}(0)}-\frac{\rho_{G}^{i}(k)}{\rho_{G}^{i}(0)}\right|.$
  • Corr metric on feature dependency: For $d>1$, we assess the feature dependency by using the $l_{1}$ norm of the difference between cross-correlation matrices and name it as Corr metric. To be specific, let $\tau^{i,j}_{r}$ and $\tau^{i,j}_{G}$ denote the correlation of the $i^{th}$ and $j^{th}$ feature of time series under real measure and synthetic measure resp. The correlation metric between the real data and synthetic data is given by $l_{1}$ norm of the difference between two correlation matrices, i.e.
    $\frac{1}{d^{2}}\sum_{i=1}^{d}\sum_{j=1}^{d}|\tau^{i,j}_{r}-\tau^{i,j}_{G}|.$
  • $R^{2}$ error for usefulnese: In order to be useful, the synthetic data should inherit the predictive characteristics of the original, meaning that the synthetic data should be just as useful as the real data when used for the same predictive purpose (i.e. train-on-synthetic, test-on-real). To measure the usefulness of the synthetic data, we follow (Esteban, Hyland, and Rätsch 2017; Yoon, Jarrett, and Van der Schaar 2019) and consider the problem of predicting next-step temporal vectors using the lagged values of time series using the real data and synthetic data. First, we train a supervised learning model on real data to predict next-step values and evaluate it in terms of $R^{2}$ (TRTR). Then we train the same supervised learning model on synthetic data and evaluate it on the real data in terms of $R^{2}$ (TSTR). The closer two $R^{2}$ are, the better the generative model is. The predictive score is then defined as the $R^{2}$ relative error. This test metric is reasonable because it demonstrates the ability of the synthetic data to be used for real applications.
Appendix E Architectures, hyperparameters, and training techniques
Image generation
Backbone architectures
We employ BigGAN (Brock, Donahue, and Simonyan 2018) and cStyleGAN2 (Karras et al. 2020b) architectures as the backbones for image generation experiments. BigGAN (Brock, Donahue, and Simonyan 2018), as a member of projection-based cGAN, is a collection of recent best practices in conditional image generation, and it is widely used due to its satisfactory generation performance on high-fidelity image synthesis. We use the BigGAN architecture with the same regularization methods like Exponential Moving Averages (EMA) (Karras et al. 2017) and Spectral Normalization (Miyato et al. 2018) have already been adopted. We adopt BigGAN’s PyTorch implementation  [^5] https://github.com/PeterouZh/Omni-GAN-PyTorch and shows the architectural details in Table 7 for completeness.
@@ Appendix E Architectures, hyperparameters, and training techniques
cStyleGAN2 (Zhao et al. 2020), is an improved and conditional version of the original StyleGAN, is a generative adversarial network (GAN) architecture designed for creating high-quality, diverse images. It addresses artifacts, enhances image quality, and has been widely used for generating realistic portraits and artwork. We adopt the code from the Github repository in (Zhao et al. 2020)  [^6] https://github.com/mit-han-lab/data-efficient-gans/tree/master. and use the default hyperparameter setting. The only difference is the minor modification we made to incorporate our MC method.
Loss functions
Since our MCGAN replaces the original generative loss in Eqn. (Conditional GANs) with the regression loss in Eqn. (3) and keeps the discriminator loss, we use two popular discriminator’s loss functions as baselines: the Hinge loss baseline and the BCE loss baseline.
Hyperparameters
For the CIFAR-10 experiment, the batch size is set to 32. We adopt the Adam optimizer in all experiments, with betas being 0.0 and 0.999. For both of the generator and discriminator, the learning rates are set to 0.0002 and the weight decay is 0.0001. The model is updated by using Exponential Moving Average starting after the first 5000 iterations. The generator is updated once every 3 times the discriminator is updated. For the CIFAR-100 experiment, we have fewer training samples for each class, so we update the discriminator 4 times per generator training as a more accurate discriminator is needed. Each experiment is conducted on one Quadro RTX 8000 GPU. All experiments are conducted using fixed and default seed settings in the code base.
For the large-scale and high-resolution experiments, our MCGAN uses the Monte Carlo sample size $M=10$ for the LSUN bedroom dataset, and $M=4$ for FFHQ256 and ImageNet64 datasets. Their detailed settings are provided in the code released.
Leaky Clamp
where $\alpha\in(0,1)$ is a small slope for values outside the range $[lb,ub]$. Just similar to the negative slope in Leaky ReLU (Maas et al. 2013), the $\alpha$ in the Leaky Clamp function is used to prevent from vanishing gradient problem. And by applying this Leaky Clamp on the discriminator output when computing the regression loss, we are able to mitigate the early collapse problem.
Data augmentation
To alleviate the overfitting and improve the generalization on the small training set, especially for CIFAR-100 where each class has scarce samples, we increase data efficiency by using the Differentiable Augmentation (DiffAug) (Zhao et al. 2020) which imposes various types of augmentations on real and fake samples (Zhao et al. 2020; Karras et al. 2020a; Zhang et al. 2019). We adopt Translation + Cutout policy as suggested in (Zhao et al. 2020). Besides, we also apply horizontal flips when loading the training dataset as in (Kang et al. 2021).
Video generation
Backbone architecture
For both the generator and discriminator, the backbone we used in conditional video generation task is called convolutional LSTM (ConvLSTM) unit proposed by (Shi et al. 2015) due to its effectiveness in video prediction tasks. For both generator and discriminator, the number of layers is 2 and the hidden dimension is specified as 64. The convolutional kernel size is set as (3,3) with padding (1,1). The activation function used is ReLU. In the model training, the generator takes in 5 past frames as the input and generates the corresponding 1-step future frame, then the real past frames and the generated future frames are concatenated along time dimension and put into the discriminator.
Hyperparameter
For moving MNIST dataset, the batch size is set to 16. And the frame size is downsampled from 64 to 32 and MC size is specified as 4 to reduce GPU memory consumption. The generator is updated every time the discriminator is updated. The Adam optimizer is adopted, with betas being 0.5 and 0.999 and learning rate being 2e-4.
Timeseries generation
Backbone architecture
The RNN model we employed in conditional timeseries generation task is built up using the AR-FNN architecture introduced in (Liao et al. 2024). The AR-FNN is defined as a composition of PReLUs, residual layers and affine transformations. Its inputs are the past $p$-lags of the $d$-dimensional process we want to generate as well as the $d$-dimensional noise vector. A formal definition can be found in (Liao et al. 2024).
Hyperparameter
For both low-dim VAR and stock datasets, the hidden dimension of AR-FNN is set as 50 with 3 number of layers. To increase model’s capacity in generating high-dim time-series, we increased the hidden dimension to 128 for $d=10,50,100$ in VAR experiment. Similar to image generation experiment, the Adam optimizer is used with betas being (0, 0.9) and learning rate being 2e-4 for both discriminator and generator. The number of total training epochs is 1000 with batch size specified as 100. The generator is updated every 4 times the discriminator is updated. For all time-series experiment, we set MC size as 1000 due to less GPU memory consumed for time-series data.
Appendix F Supplementary numerical results
In this section, we present the supplementary numerical results on both image generation and timeseries generation.
Image generation
Sensitivity analysis of MC sample size
In our experiments, we use a Monte Carlo sample size ($M$) of 10 for BigGAN and 4 for StyleGAN2 as outlined. We’ve conducted a sensitivity analysis using BigGAN on CIFAR-10 datasets. The results can be found in Table 8.
w/o MC $M$=5 $M$=10 $M$=15 $M$=20 $M$=25
FID $\downarrow$ 4.34 3.73 3.55 3.61 4.99 4.50
IS $\uparrow$ 9.41 9.69 9.96 10.07 10.32 10.16
Time (s) 976 1560 2171 2868 3498
Table 8: Sensitivity analysis of MC sample size ($M$) based on BigGAN trained on CIFAR-10 dataset. The training time (seconds) is computed for 5k iterations.
A sweet spot around $M=10$ is observed with the best FID and a good IS. A larger MC size of fake samples increases the variability in the IS score but causes the generated distribution to diverge from the real data. The optimal size may vary for different datasets, indicating the need (limitation) for fine-tuning. Future work could explore its adaptive strategies.
In Table 8, we also report the training time comparison between the baseline and our MCGAN with varying $M$ for our BigGAN CIFAR-10 experiments. In general, the training time of MCGAN is approximately linear w.r.t Monte Carlo size $M$ for the fixed number of epochs. Since a moderate $M$ achieves satisfactory results and MCGAN often converges faster, the training time remains manageable compared to the baseline.
Failure case study on FFHQ256
To evaluate the quality of generated samples, we compared the baseline model (StyleGAN2-ada) with our proposed method (StyleGAN2-ada+MC). We generated 480 samples from each model for a qualitative analysis. Our observations reveal that only the baseline model produces samples with missing facial components, as illustrated in Figure 7(a). This suggests that our method captures facial structures more effectively. Furthermore, generating coherent faces under occlusion is challenging. As shown in Figures 7(b) and 7(c) , our method produces more realistic facial structures behind the microphone, which we attribute to the strong supervision provided by our innovative loss function.
(a) Samples with missing facial features generated by StyleGAN2-ada;
(b) Samples with mic generated by StyleGAN2-ada;
(c) Samples with mic generated by StyleGAN2-ada + MC.
Figure 7: Failure cases generated by baseline and our model
Timeseries generation
In this section, we present our numerical results on conditional timeseries generation .
VAR dataset
For VAR dataset, the evaluation metrics are give in Tables 9, 10 and 11. From these tables, we can see that our MCGAN has considerable improvement over RCGAN across different parameter settings and dimensions. Also, the long-term ACF shown in Figure 8 illustrates that MCGAN can better capture the temporal dependence than RCGAN.
Table 9: Numerical results of $\operatorname{VAR}(1)$ for $d=1$
Temporal Correlations
Settings $\phi=0.2$ $\phi=0.5$ $\phi=0.8$
Metric on marginal distribution
SigCWGAN 0.00522 0.00610 0.00381
MCGAN 0.00402 0.00501 0.00384
TimeGAN 0.0259 0.02735 0.01691
RCGAN 0.00443 0.00683 0.00464
GMMN 0.00678 0.00659 0.00554
Absolute difference of lag-1 autocorrelation
SigCWGAN 0.00947 0.01464 0.00182
MCGAN 0.00648 0.02047 0.00324
TimeGAN 0.04269 0.04526 0.01651
RCGAN 0.00266 0.01943 0.00531
GMMN 0.01232 0.00106 0.00618
Relative $R^{2}$ error (%)
SigCWGAN 0.45011 0.12953 0.00654
MCGAN 0.15403 0.39642 0.03417
TimeGAN 7.44523 2.12036 1.38983
RCGAN 2.16534 0.93133 0.19214
GMMN 0.34882 1.36565 2.10632
Sig-$W_{1}$ distance
SigCWGAN 0.69598 1.09869 2.34807
MCGAN 0.69529 1.10365 2.35118
TimeGAN 0.71696 1.12885 2.37692
RCGAN 0.69653 1.0995 2.35203
GMMN 0.70083 1.10592 2.3526
Table 10: Numerical results of $\operatorname{VAR}(1)$ for $d=2$
Temporal Correlations (fixing $\sigma=0.8$) Feature Correlations (fixing $\phi=0.8$)
Settings $\phi=0.2$ $\phi=0.5$ $\phi=0.8$ $\sigma=0.2$ $\sigma=0.5$ $\sigma=0.8$
Metric on marginal distribution
SigCWGAN 0.01177 0.00537 0.00365 0.00383 0.00277 0.00365
MCGAN 0.00384 0.00651 0.00538 0.00457 0.00502 0.00538
TimeGAN 0.02059 0.02187 0.01113 0.00933 0.01099 0.01113
RCGAN 0.00613 0.00706 0.00466 0.00607 0.00886 0.00466
GMMN 0.00861 0.00912 0.00601 0.00474 0.00476 0.00601
Absolute difference of lag-1 autocorrelation
SigCWGAN 0.00658 0.00248 0.00419 0.00353 0.00555 0.00419
MCGAN 0.02137 0.04051 0.00716 0.00438 0.00840 0.00716
TimeGAN 0.04433 0.04567 0.00822 0.02446 0.00442 0.00822
RCGAN 0.01857 0.04249 0.03218 0.01227 0.03571 0.03218
GMMN 0.00699 0.02081 0.04263 0.08085 0.05893 0.04263
$L_{1}$-norm of real and generated cross-correlation matrices
SigCWGAN 0.00804 0.01113 0.01122 0.00476 0.01198 0.01122
MCGAN 0.02653 0.01502 0.01149 0.01381 0.03842 0.01149
TimeGAN 0.08622 0.07002 0.07494 0.07455 0.04685 0.07494
RCGAN 0.01200 0.02846 0.03460 0.08187 0.03317 0.03460
GMMN 0.00745 0.00565 0.02705 0.00973 0.00917 0.02705
Relative $R^{2}$ error (%).
SigCWGAN 1.24036 0.09027 0.01252 0.01381 0.01248 0.01252
MCGAN 10.24923 1.61850 0.42136 0.26449 0.35319 0.42136
TimeGAN 40.1273 4.92783 1.21018 1.05100 0.89636 1.21018
RCGAN 18.33682 4.31191 1.39435 3.94201 1.58417 1.39435
GMMN 35.25094 15.76457 6.56956 12.42385 9.88914 6.56956
Sig-$W_{1}$ distance
SigCWGAN 1.92823 2.42590 3.60068 3.02208 3.23497 3.60068
MCGAN 1.93087 2.42466 3.61617 3.02879 3.2390 3.61617
TimeGAN 1.98070 2.47622 3.63571 3.04472 3.26746 3.63571
RCGAN 1.93333 2.43379 3.61464 3.03564 3.21083 3.61464
GMMN 1.94517 2.43949 3.60922 3.02910 3.23898 3.60922
Table 11: Numerical results of $\operatorname{VAR}(1)$ for $d=3$
Temporal Correlations (fixing $\sigma=0.8$) Feature Correlations (fixing $\phi=0.8$)
Settings $\phi=0.2$ $\phi=0.5$ $\phi=0.8$ $\sigma=0.2$ $\sigma=0.5$ $\sigma=0.8$
Metric on marginal distribution
SigCWGAN 0.01463 0.01240 0.00477 0.00423 0.00452 0.00477
MCGAN 0.00476 0.00436 0.00596 0.00715 0.00661 0.00596
TimeGAN 0.02359 0.02096 0.00886 0.01054 0.00915 0.00886
RCGAN 0.01068 0.00634 0.00577 0.00836 0.00597 0.00577
GMMN 0.01001 0.01024 0.00987 0.01427 0.01323 0.00987
Absolute difference of lag-1 autocorrelation
SigCWGAN 0.00570 0.00508 0.00131 0.00330 0.00172 0.00131
MCGAN 0.00684 0.01805 0.0199 0.00947 0.00529 0.0199
TimeGAN 0.04601 0.09309 0.01643 0.03144 0.04736 0.01643
RCGAN 0.05663 0.04925 0.02041 0.01894 0.01863 0.02041
GMMN 0.04041 0.06024 0.08998 0.10196 0.13395 0.08998
$L_{1}$-norm of real and generated cross-correlation matrices
SigCWGAN 0.01214 0.01311 0.00317 0.01715 0.02862 0.00317
MCGAN 0.04076 0.03819 0.03659 0.04631 0.08001 0.03659
TimeGAN 0.20056 0.43239 0.15509 0.09314 0.09228 0.15509
RCGAN 0.24082 0.16809 0.09657 0.16257 0.11514 0.09657
GMMN 0.09850 0.12638 0.20142 0.3096 0.37507 0.20142
Relative $R^{2}$ error (%).
SigCWGAN 1.393190 0.34009 0.07690 0.05323 0.03498 0.07690
MCGAN 14.63272 1.62564 0.56412 0.32549 0.42388 0.56412
TimeGAN 36.71498 8.94899 2.38110 2.61944 3.80723 2.38110
RCGAN 70.69909 16.50512 2.83140 1.44543 2.79532 2.83140
GMMN 152.87792 38.97992 17.94085 25.12542 26.93346 17.94085
Sig-$W_{1}$ distance
SigCWGAN 11.57390 17.66105 30.46722 30.75008 25.24824 30.46722
MCGAN 11.57797 17.69298 30.48571 30.73360 25.22319 30.48571
TimeGAN 11.88320 18.09083 30.70047 30.86857 25.36035 30.70047
RCGAN 11.53368 17.72101 30.40070 30.74105 25.30295 30.40070
GMMN 11.61313 17.73444 30.59544 30.79028 25.38754 30.59544
Figure 8: ACF plot for each channel on the 3-dimensional VAR(1) dataset with autocorrelation coefficient $\phi=0.8$ and co-variance parameter $\sigma=0.8$. Here $x$-axis represents the lag value ( with a maximum lag equal to 100) and the $y$-axis represents the corresponding auto-correlation. The length of the real/generated time series used to compute the ACF is 1000. The number in the bracket under each model is the sum of the absolute difference between the correlation coefficients computed from real (dashed line) and generated (solid line) samples.
Stocks dataset
For stock dataset, we generate both log return and log volatility process of S&P 500 and DJI . The estimated cross-corelation matrices are presented in Figure 10, we can see that the generated cross-corelation matrices by MCGAN are closer to the ground truth comparing with that of RCGAN, indicating the effectiveness of our MC method in capturing cross dependency.
Figure 10: Comparison of real and synthetic cross-correlation matrices for SPX/ SPX and DJI data. On the far left the real cross-correlation matrix from SPX/ SPX and DJI log-return and log-volatility data is shown. $x$/$y$-axis represents the feature dimension while the color of the $(i,j)^{th}$ block represents the correlation of $X_{t}^{(i)}$ and $X_{t}^{(j)}$. Observe that the historical correlation between log returns and log volatility is negative, indicating the presence of leverage effects, i.e. when log returns are negative, log volatility is high.
Synthetic 2D dataset
We conducted experiments using 2D grid-like synthetic data with 25 Gaussian modes and 5,000 generated samples. Our MCGAN registered all 25 modes with Total Variation TV=$14.64\pm 5.50$, outperforming vanilla GAN (17 modes, TV=$35.23\pm 2.02$) and LSGAN (20 modes, TV=$29.72\pm 6.40$), indicating our ability to alleviate mode collapse. Increasing MC size to 50/100 reduced TV further to $6.99\pm 4.67$/$3.54\pm 2.17$ as shown in Table 12. Visualizations are given in Figure 11.
Table 12: Results on 2D synthetic data. Test metrics are computed using 5000 generated samples for 10 different seeds. The label MC=$n$ indicates that MCGAN is used with a Monte Carlo sample size of $M=n$.
Method # Registered Modes # Registered Points Total Variation
GAN $17.4\pm 3.1$ $4511.8\pm 63.86$ $35.23\pm 2.02$
LSGAN $20.4\pm 1.2$ $4464.2\pm 182.85$ $29.72\pm 6.40$
MC =10 $25\pm 0.0$ $4659.4\pm 75.72$ $14.64\pm 5.50$
MC=50 $25\pm 0.0$ $\mathbf{4807.8\pm 21.07}$ $6.99\pm 4.68$
MC=100 $\mathbf{25\pm 0.0}$ $4800.4\pm 59.85$ $\mathbf{3.54\pm 2.17}$
(a) Vanilla GAN (TV=38.62)
(b) LSGAN (TV=24.97)
(c) MC=10 (TV=14.40)
(d) MC=100 (TV=2.84)
Figure 11: Example of generated samples by different methods; Red points are 5000 real samples with 0.01 standard deviation; Blue points are 5000 generated samples; Dash lines illustrate gradients of discriminator on each point.
Appendix G Generated Samples
In this section, we show some generated samples generated by our MCGAN models including images, time-series.
Generated CIFAR-10 samples by BigGAN backbone
Here we present the samples generated by BigGAN backbone in Figure 3. We can see that only a few generated figures are misclassified, showcasing the ability of MCGAN in genrating high-fidility samples.
Generated CIFAR-10 samples by StyleGAN2 backbone
Here we present the samples generated by StyleGAN2 backbone in Figure 9. Comparing with Figure 3, fewer generated samples are misclassified due to the employment of a much stronger backbone.
Generated FFHQ256 samples by StyleGAN2 backbone
In this subsection, we present some FFHQ256 samples generated by StylGAN2 trained via our MC methods in Figure 11. These 64 images are randomly picked out of 480 generated samples. We can observe from Figure 11 human faces with different skins, ages, angles, lighting and accessories, showcasing that StyleGAN2 trained via our MC method has the ability to generate realistic, diversified, and high-resolution human face images.
Figure 11: FFHQ256 samples generated by the StyleGAN2 backbone trained via our MC method with FID $3.77\pm 0.04$.
Generated ImageNet64 samples by StyleGAN2 backbone
In this subsection, we present 100 samples generated by cStylGAN2 backbone trained via our MC methods in Figure 12. Due to the large scale of ImageNet64 dataset, it is a rather challenging conditional generation task.
Figure 12: 100 ImageNet64 samples generated by the cStyleGAN2 backbone trained via our MC method with FID $16.76\pm 0.08$.
Generated LSUN bedroom samples by StyleGAN2 backbone
In this subsection, we present 100 samples generated by cStylGAN2 backbone trained via our MC methods in Figure 13.
Figure 13: 100 LSUN bedroom samples generated by the cStyleGAN2 backbone trained via our MC method with FID $2.77\pm 0.03$.
Generated stock data samples by RNN backbone
In this subsection, we present generated SPX log-return paths of each model in Figure 14. We can see that visually the log-return path generated by MCGAN shows volatility clustering property and looks closer to the historical path than that of RCGAN. We also provide comparison of marginal distributions of ground truth and generated paths of MCGAN in Figure 15. We can see that the generated histogram is very close to the historical one in terms of mean, standard deviation, skewness and kurtosis.
(a) Real
(b) MCGAN
(c) SigCWGAN
(d) TimeGAN
(e) RCGAN
(f) GMMN
Figure 14: Example paths of SPX log returns generated by each model. Since the path of DJI log returns is similar to that of SPX, there is no need to make another plot for DJI.
Figure 15: Comparison of the marginal distributions of the paths generated by MCGAN and the SPX and DJI data.